OpenStudy (anonymous):

In the function f(x) = 6x how does an increase of 1 in the independent variable x affect f(x)? f(x) will have a decrease of 6x - 6 f(x) will have an increase of 6 f(x) will have an increase of 6x + 6 f(x) will have a decrease of 6

OpenStudy (anonymous):

Increase of 6: f(1)=6(1)=6

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

thanks:))

OpenStudy (anonymous):

Think about it... 6*1 = 6, 6*2 = 12 (x increased by one and therefore, the function output increased by 6).